With a tailwind, a small aircraft can fly 900 miles in 5 hours. Against the same wind, the plane can fly the same distance in 6 hours. Find the average wind speed and the average air speed of the plane.
Answer by Alan3354(30937)

With a tailwind, a small aircraft can fly 900 miles in 5 hours. Against the same wind, the plane can fly the same distance in 6 hours. Find the average wind speed and the average air speed of the plane.
---------------
900/5 = 180 mph
900/6 = 150 mph
The plane's airspeed is the average = (180 + 150)/2 = 165 mph
The windspeed is the difference = 15 mph